OpenStudy (anonymous):

A skater traveling at 7 m/s accelerates at the rate of .80m/s/s for an interval of 2.0sec. What is the final velocity? ***thank you!! :)

OpenStudy (phi):

any ideas?

OpenStudy (anonymous):

f=m*a ? not sure :/

OpenStudy (phi):

v= a*t

OpenStudy (michele_laino):

what is the formula which gives the velocity of an object in acceleration?

OpenStudy (michele_laino):

@iheartfood

OpenStudy (anonymous):

ohh okay! so 7*2? or is it 7*0.8 ?

OpenStudy (anonymous):

for v=a*t?

OpenStudy (michele_laino):

hint, your formula is: \[v(t)=v _{0}+a*t\] where v_0 is the initial velocity @iheartfood

OpenStudy (phi):

you should try to match units. also, strictly speaking, v=at is the change in velocity so we add that to the original velocity (if any)

OpenStudy (phi):

if we look only at the units velocity is m/s acceleration is m/s^2 notice m/s^2 * s (i.e. a * t) gives units of m/s (velocity)

OpenStudy (anonymous):

ohh so i would want to do 7 m/s * 0.8 m/s ?

OpenStudy (anonymous):

and what would v0 be? :/

OpenStudy (michele_laino):

v_0=7m/sec

OpenStudy (phi):

that would give m^2/s^2 and that is not velocity

OpenStudy (anonymous):

ohhhh wait so what would it look like to substitute? 7 + 0.8 * 2 ? so 7+1.6 = 8.6 ?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

oh! so that would be the final velocity? like this? 8.6 m/s^2 ?

OpenStudy (phi):

yes, but with these problems always put in the units. (It is safer that way, and helps keep things clear)

OpenStudy (anonymous):

ahh okie! :) and it would be m/s^2 for the final velocity? :O

OpenStudy (michele_laino):

no, unit of measure of velocity is m/sec

OpenStudy (phi):

and you should be m/s as the final units (velocity is distance divided by time, e.g. meter/seconds

OpenStudy (anonymous):

ohh so it's 8.6 m/s ? :)

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

okay! awesome!! thank you both!! @phi @Michele_Laino :)

OpenStudy (michele_laino):

thank you! @phi @iheartfood :)

ganeshie8 (ganeshie8):

distance : m velocity = m/s acceleration : (m/s)/s jerk : ((m/s)/s)/s ...

OpenStudy (michele_laino):

sorry, what is "jerk" please ? @ganeshie8

ganeshie8 (ganeshie8):

http://en.wikipedia.org/wiki/Jerk_%28physics%29

OpenStudy (michele_laino):

ok! thanks! for your answer @ganeshie8

ganeshie8 (ganeshie8):

it is used very rarely i guess..

OpenStudy (michele_laino):

yes! that's true! It is mye first time that I see it

OpenStudy (michele_laino):

oops...it is my first time...